# Unified Supervision for Walmart’s Sponsored Search Retrieval via Joint Semantic Relevance and Behavioral Engagement Modeling

### 10 Apr 2026 v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne, VIC, Australia.booktitle: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’26), July 20–24, 2026, Melbourne, VIC, Australiaisbn: 979-8-4007-2599-9/2026/07doi: 10.1145/XXXXXX.XXXXXXccs: Applied computing E-commerce infrastructureccs: Information systems Novelty in information retrieval
@@ 3.1. Graded relevance label
We annotate each QIP with an ordinal relevance rating $Rel(q,i)\in\{0,1,2,3,4\}$, corresponding to 0 (Embarrassing), 11 (Bad), 22 (Okay), 33 (Good), and 44 (Excellent). This scale represents the degree of intent fulfillment, ranging from fundamental category mismatches (0) to precise, ideal matches (44) that satisfy all query constraints. Ratings are produced by a combination of available human annotations and a cascade of relevance models (fine-tuned on internal data) evaluated sequentially from least to most computationally expensive (Gemma-1B, Gemma-2B, and a LLaMA-3 8B model). Each stage outputs a 5-class predictive distribution, and a prediction is accepted early when the confidence score exceeds a stage-specific threshold. For pairs not accepted early, we run all stages, take the majority label, and break ties using the final-stage prediction. We map the relevance rating to a normalized score in $[0,1]$.
@@ 3.2. Multi channel retrieval prior score
$3.$ User engagement signals
@@ 3.4. Unified supervision and QIP scoring
Positive QIP relevance score: We combine the normalized relevance score defined in Eqn. 1 with the aggregated rank-prior score $\pi(q,i)$ and multi-channel consensus score $\mathcal{C}(q,i)$ defined in section 3.2. We define the unified relevance-rank score as:
@@
3.5. Retrieval Model Training
Figure 3. Green text highlights phrases that align with user intent, while red text shows failure cases. Unified supervision (Variation-2) retrieves relevant and popular(top 1% engaged) items for each query compared to production system (Control), while relevance only scoring (Variation-1) retrieves relevant but unpopular(bottom 10% engaged) items.
4. Evaluation
4.1. Evaluation Protocol
Offline evaluation
@@ 4.2. Quantitative Results
Figure 2 measures the percentage of highly-engaged items is present in the retriever’s Top-25 list. We first convert engagement into a within-query percentile (so the 90th percentile means “top 10% most-engaged items for that query”), then measure what fraction of the Top-25 exceeds each cutoff. The x-axis reports that fraction as a percentage, while the y-axis represents the engagement percentile threshold. Across all thresholds, the engagement supervised model retrieves a higher share of high-engagement items than the relevance-only supervised model. For example, at the 50th percentile threshold, the share increases by 14.3 $p$, and at the 90th percentile threshold it increases by 7.0 pp.
@@ 5. Discussion
Md Omar Faruk Rokon is a Staff Data Scientist at Walmart specializing in applied machine learning for Information Retrieval. He earned his PhD in Computer Science in 2022, with a research focus on embedding and representation learning. His current work leverages Transformer architectures and Large Language Models to enhance semantic understanding and search relevance. Additionally, he develops rigorous search evaluation methodologies to drive improvements in e-commerce systems

### 15 Apr 2026 v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne, VIC, Australiabooktitle: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’26), July 20–24, 2026, Melbourne, VIC, Australiadoi: 10.1145/3805712.3808420isbn: 979-8-4007-2599-9/2026/07ccs: Applied computing E-commerce infrastructureccs: Information systems Novelty in information retrieval
@@ 3.1. Graded relevance label
We annotate each QIP with an ordinal relevance rating $Rel(q,i)\in\{0,1,2,3,4\}$, corresponding to 0 (Embarrassing), $1$ (Bad), $2$ (Okay), $3$ (Good), and $4$ (Excellent). This scale represents the degree of intent fulfillment, ranging from fundamental category mismatches (0) to precise, ideal matches ($4$) that satisfy all query constraints. Ratings are produced by a combination of available human annotations and a cascade of relevance models (fine-tuned on internal data) evaluated sequentially from least to most computationally expensive (Gemma-1B, Gemma-2B, and a LLaMA-3 8B model). Each stage outputs a 5-class predictive distribution, and a prediction is accepted early when the confidence score exceeds a stage-specific threshold. For pairs not accepted early, we run all stages, take the majority label, and break ties using the final-stage prediction. We map the relevance rating to a normalized score in $[0,1]$.
@@ 3.2. Multi channel retrieval prior score
3.3. User engagement signals
@@ 3.4. Unified supervision and QIP scoring
Positive QIP relevance score: We combine the normalized relevance score defined in Eqn. 1 with the aggregated rank-prior score $\pi(q,i)$ and multi-channel consensus score $\mathcal{C}(q,i)$ defined in section 3.2.
Figure 2. Qualitative results: Green text highlights phrases that align with user intent, while red text shows failure cases. Unified supervision (Variation-2) retrieves relevant and popular(top 1% engaged) items for each query compared to production system (Control), while relevance only scoring (Variation-1) retrieves relevant but unpopular(bottom 10% engaged) items.
We define the unified relevance-rank score as:
@@
3.5. Retrieval Model Training
4. Evaluation
4.1. Evaluation Protocol
Offline evaluation
@@ 4.2. Quantitative Results
Figure 3 measures the percentage of highly-engaged items is present in the retriever’s Top-25 list. We first convert engagement into a within-query percentile (so the 90th percentile means “top 10% most-engaged items for that query”), then measure what fraction of the Top-25 exceeds each cutoff. The x-axis reports that fraction as a percentage, while the y-axis represents the engagement percentile threshold. Across all thresholds, the engagement supervised model retrieves a higher share of high-engagement items than the relevance-only supervised model. For example, at the 50th percentile threshold, the share increases by 14.3 pp, and at the 90th percentile threshold it increases by 7.0 pp.
@@ 5. Discussion
Md Omar Faruk Rokon is a Staff Data Scientist at Walmart specializing in applied machine learning for Information Retrieval. He earned his Ph.D. in Computer Science in 2022, with a research focus on embedding and representation learning. His current work leverages Transformer architectures and Large Language Models to enhance semantic understanding and search relevance. Additionally, he develops rigorous search evaluation methodologies to drive improvements in e-commerce systems